### v2

# MirrorBench: A Benchmark to Evaluate Conversational User-Proxy Agents for Human-Likeness

Abstract.
Large language models (LLMs) are increasingly used as human simulators, both for evaluating conversational systems and for generating fine-tuning data. However, naive “act-as-a-user” prompting often yields verbose, unrealistic utterances, motivating principled evaluation of user proxy agents. We present MirrorBench, a reproducible and extensible benchmarking framework that evaluates user proxies solely on their ability to produce human-like user utterances across diverse conversational regimes, explicitly decoupled from downstream task success. MirrorBench combines three lexical-diversity metrics (MATTR, Yule’s $K$, and HD-D) with three LLM-judge-based metrics (GTEval, Pairwise Indistinguishability, and Rubric-and-Reason), and contextualizes judge scores using Human–Human and Proxy–Proxy calibration controls. Across four public datasets, MirrorBench yields variance-aware comparisons and reveals systematic gaps between user proxies and real human users. The framework is open source [^1] https://github.com/SAP/mirrorbench and includes a command-line interface for running and managing user-proxy benchmarking experiments.
Large Language Models (LLMs), User-proxy agents, Human-likeness benchmarking, LLM-as-a-judge, Judge sensitivity, Human–LLM correlation, Lexical diversity, MATTR, HD-D, Yule’s K, Pairwise indistinguishability, GTEval, Rubric-And-Reason, ChatbotArena, ClariQ, QULAC, OASST1
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Preprint Under Review; Feb 2025; isbn: 978-1-4503-XXXX-X/2026/08ccs: Computing methodologies Intelligent agentsccs: Computing methodologies Discourse, dialogue and pragmaticsccs: Computing methodologies Lexical semanticsccs: Computing methodologies Natural language generationccs: Computing methodologies Simulation evaluation
1. Introduction
Deploying and improving conversational AI hinges on evaluation under realistic user interactions. Expert human-annotated utterances, while authoritative, is costly and difficult to scale for systematic evaluation. At the same time, the instruction-tuning of models for production systems demands large volumes of authentic user–assistant exchanges. To meet both needs, practitioners increasingly rely on user proxy agents, namely LLMs prompted to emulate specified human personas, to synthesize user utterances for automated testing and post-training at scale.
In practice, user proxies are now used far beyond toy role-play: they drive regression testing for assistant updates (Wang et al., 2025b), evaluate tool-use and grounding behaviors (Zhang et al., 2025), generate synthetic dialogues for post-training (Naous et al., 2025; Hathidara et al., 2025), and enable stress tests that are infeasible with human participants (Lynch et al., 2025; Zhang and Yang, 2025). However, naive “act-as-a-user” prompting often yields verbose, overly cooperative utterances that diverge from real user behavior (Zhou et al., 2024), indicating that unrealistic proxies can bias downstream evaluations and introduce distribution shift when used for data generation. This motivates a central question: How can we rigorously measure the human-likeness of a user proxy, independently of downstream task success? Here, we define a user as a human interlocutor participating in a specific conversational or task-oriented dialogue.
We introduce MirrorBench, a reproducible and extensible framework for variance-aware user-proxy benchmarking on human-likeness metrics using multiple human reference datasets across diverse conversational settings. MirrorBench evaluates proxies solely on human-likeness and deliberately separates this from task completion. We define the human-likeness of a user proxy as the extent to which its generated user utterances match real human user utterances along two complementary axes: (i) lexical similarity, operationalized via human-anchored lexical-diversity statistics (e.g., MATTR, HD-D, Yule’s $K$), and (ii) behavioral realism, operationalized via LLM-judge evaluations that assess whether the proxy’s user turns sound like a real user in context (e.g., naturalness, tone, appropriateness).
Human-likeness is inherently scenario-dependent: what sounds realistic in open-domain chat may be unrealistic in information-seeking or clarification-heavy interactions. Accordingly, MirrorBench does not aim to capture an unconstrained notion of “general” human behavior. Instead, we evaluate utterance-level realism conditioned on the same task context, asking whether a proxy produces user turns that resemble those of real users in that regime. This framing explicitly decouples user-proxy quality from assistant task success, and treats human-likeness as a measurable property of the user side of the interaction under controlled conditions.
To evaluate the realism of user–proxy agents against real-world dialogues, MirrorBench packages four open-source corpora preprocessed for user-proxy evaluation: QULAC (Aliannejadi et al., 2019), ClariQ (Aliannejadi et al., 2021), OASST1 (Köpf et al., 2023), and ChatbotArena (Zheng et al., 2023). To quantify alignment in lexical diversity with human users, we implement Moving-Average Type-Token Ratio (MATTR) (Covington and McFall, 2010), Yule’s $K$ (Tanaka-Ishii and Aihara, 2015), and a Hypergeometric Distribution Diversity (HD-D) (Carmona et al., 2019). To assess behavioral realism, we employ a family of LLM-judge metrics, including GTEval (Zhu et al., 2025), Pairwise Indistinguishability (PI) (Zheng et al., 2023), and Rubric-and-Reason (RNR) (Liu et al., 2023b). Taken together, these metrics quantify the realism of the user side of the dialogue, explicitly decoupled from assistant capabilities and task success.
Our benchmark reveals that even strong LLMs exhibit systematic gaps from real users, and that these gaps vary substantially across conversational regimes. Across datasets, we observe a recurring realism-diversity tension: proxies that score highly on judge-based behavioral realism can still under-match human lexical/distributional characteristics in clarification-centric settings, motivating the use of complementary metric families. We also find that absolute judge scores and, in some cases, fine-grained orderings can shift with the judge choice, reinforcing the importance of transparent multi-judge reporting and calibration controls when comparing closely matched proxies.
In summary, our contributions are as follows: (1) a benchmark protocol for evaluating user-proxy agents for human-likeness with human-anchored normalization and calibrated judge metrics; (2) six metrics spanning lexical-diversity (MATTR, HD-D, Yule’s $K$) and judge-based realism (GTEval, Pairwise Indistinguishability, Rubric-and-Reason); (3) a unified preprocessing of four open conversational datasets for user-proxy benchmarking; and (4) an empirical study across five user-proxy LLMs, yielding actionable insights on realism-diversity tensions and judge sensitivity.
2. Related Work
User simulation and user proxies.
Simulating users to train and evaluate dialogue agents has a long history in NLP (Balog and Zhai, 2023). Early user simulators were typically goal-driven or rule-based, using predefined agendas over dialogue acts to approximate human behavior (Schatzmann and Young, 2009). More recently, LLM-based user proxies have enabled open-ended interactions that better match real conversational variability, but they also introduce failure modes such as verbosity, role drift, and inconsistent intent adherence. For instance, (Wang et al., 2025a) proposes a User Simulator with Implicit Profiles (USP) that infers latent user traits to improve realism beyond naive role-play prompting. LLM user proxies are increasingly used as evaluation drivers in interactive testbeds (e.g., tool-use and multi-turn assistant evaluation) (Lu et al., 2025; Ahmad et al., 2025; Hathidara et al., 2025), and related systems generate synthetic human–assistant conversations from seed dialogues (Zhu et al., 2025). In parallel, broader agent evaluation and simulation frameworks (e.g., AgentBench (Liu et al., 2025) and AgentScope (Pan et al., 2024)) evaluate agents in environments or support large-scale multi-agent rollouts, but do not isolate and benchmark the human-likeness of the user role as a first-class object. In contrast, MirrorBench evaluates user proxies directly by comparing their produced user utterances against real-user distributions across multiple conversational regimes, combining judge-based behavioral realism with distributional/lexical realism.
LLMs as judges.
LLM-as-a-judge evaluation prompts a strong model to rate or compare dialogue quality, reducing reliance on reference answers or large-scale human annotation (Liu et al., 2023a; Li et al., 2023). Systems such as MT-Bench/ChatbotArena (Zheng et al., 2023), G-Eval (Liu et al., 2023a), and rubric-driven evaluators such as Prometheus (Kim et al., 2024) have enabled scalable, flexible assessments by operationalizing custom criteria in natural language prompts. However, most LLM-judge setups focus on evaluating assistant outputs, whereas our setting repurposes judge-based evaluation to assess the user side and pairs it with explicit calibration controls (HH/PP) and complementary lexical/distributional metrics to better contextualize score magnitudes across datasets and model families.
3. MirrorBench
This section describes the MirrorBench evaluation methodology for evaluating user-proxy LLMs across diverse conversational datasets. Figure 1 summarizes the MirrorBench pipeline. We start from diverse conversational datasets and synthesize user goals to drive controlled, goal-conditioned rollouts between a user-proxy LLM and an assistant LLM under explicit role separation. The resulting transcripts are scored with a combined evaluation suite of judge-based behavioral realism metrics and lexical/distributional metrics, with Human–Human (HH) and Proxy–Proxy (PP) controls used to contextualize judge scores. Finally, we run robustness analyses, swapping assistants or judges and repeating runs across seeds, to quantify sensitivity and variance before reporting benchmark outputs such as leaderboards, reliability checks, and operational profiles. Detailed system architecture is described in Appendix A.
Figure 1. MirrorBench benchmark pipeline, spanning dataset preparation, synthetic rollouts, the evaluation suite, and an overlay of robustness/sensitivity analysis.
3.1. Benchmark Protocol
We formalize evaluation of a user–proxy agent $\theta_{u}$ on a domain-agnostic dialogue dataset $D^{\mathrm{ref}}$ with respect to a set of human-likeness metrics $M$. Let
$D^{\mathrm{ref}}\;=\;\{\,d^{\mathrm{ref}}_{j}\in\mathcal{D}\,\}_{j=1}^{n}$
denote a corpus of reference, human-grounded dialogues (where $\mathcal{D}$ is the set of finite dialogues). Each dialogue is a sequence of user/assistant turns,
$d^{\mathrm{ref}}_{j}\;=\;\bigl[(u^{\mathrm{ref}}_{j,1},a^{\mathrm{ref}}_{j,1}),\ldots,(u^{\mathrm{ref}}_{j,L_{j}},a^{\mathrm{ref}}_{j,L_{j}})\bigr],$
with variable length $L_{j}\!\geq\!1$. Here $u^{\mathrm{ref}}_{j,t}$ is a human user utterance at turn $t$, and $a^{\mathrm{ref}}_{j,t}$ is the corresponding assistant reply (curated by a human expert or generated by an LLM in the source dataset).
Goal specification.
In some datasets, each reference dialogue $d^{\mathrm{ref}}_{j}$ includes an explicit user-goal annotation $g_{j}\in\mathcal{G}$. When no goal is provided, we derive one using an auxiliary goal generator $\theta_{\mathrm{goal}}:\mathcal{D}\to\mathcal{G}$ applied to $d^{\mathrm{ref}}_{j}$:
$g_{j}\;=\;\begin{cases}\text{given goal attached to }d^{\mathrm{ref}}_{j},&\text{if available},\\[2.0pt] \theta_{\mathrm{goal}}\!\bigl(d^{\mathrm{ref}}_{j}\bigr),&\text{otherwise}.\end{cases}$
The goal specification is used to condition $\theta_{u}$ when synthesizing user utterances; if $D^{\mathrm{ref}}$ already provides such annotations, the goal-inference step is omitted.
Synthetic rollouts.
To compare a user–proxy agent’s behavior against human references, we synthesize proxy–assistant dialogues by rolling out the user–proxy $\theta_{u}$ against an assistant model $\theta_{a}$. For each reference dialogue $d^{\mathrm{ref}}_{j}$ with user goal $g_{j}$, we condition $\theta_{u}$ on $g_{j}$ and the accumulated dialogue history to produce user turns $\hat{u}_{j,t}$, while $\theta_{a}$ produces contextual assistant replies $\hat{a}_{j,t}$. Here, we condition $\theta_{a}$ on $d^{\mathrm{ref}}_{j}$ such that the synthetic rollout trajectory follows the same path as real one. This yields a synthetic transcript
$\hat{d}_{j}\;=\;\bigl[(\hat{u}_{j,1},\hat{a}_{j,1}),\ldots,(\hat{u}_{j,L_{j}},\hat{a}_{j,L_{j}})\bigr],$
which we compare with the corresponding reference dialogue $d^{\mathrm{ref}}_{j}$ to score human-likeness of the user-proxy. Importantly, our evaluation targets the user side $\{\hat{u}_{j,t}\}_{t=1}^{L_{j}}$ only: we consider the behavior, tone, and style of the proxy’s utterances, and do not score assistant response content.
Task formalization.
Let $T$ denote the fully specified, realism-evaluation task for the human-likeness of a user-proxy agent $\theta_{u}$ over the reference corpus $D^{\mathrm{ref}}$. The task encapsulates all configuration required for execution and metric aggregation, ensuring that the evaluation process is reproducible and self-contained.
Inputs. The inputs to $T$ include a reference dataset $D^{\mathrm{ref}}$, an optional goal generator $\theta_{\mathrm{goal}}$ (used when no explicit goal is available), the user-proxy agent under test $\theta_{u}$, an assistant model $\theta_{a}$ used for synthetic rollouts, and a set of human-likeness metrics $M$.
Outputs. The evaluation can be expressed as a composite executable function
$R_{\theta_{u}}\;=\;\Psi_{T}\!\bigl(\theta_{u},\,D^{\mathrm{ref}},\,M;\,\theta_{a},\,\theta_{\mathrm{goal}}\bigr),$
where $\Psi_{T}$ denotes the instantiated evaluation pipeline. This pipeline performs rollout generation by simulating synthetic dialogues $\hat{D}=\{\hat{d}_{j}\}_{j=1}^{n}$ between $\theta_{u}$ and $\theta_{a}$ under the same dialogue structures and goals as $D^{\mathrm{ref}}$. It then applies the metric suite $M$ to each dialogue pair $(\hat{d}_{j},d^{\mathrm{ref}}_{j})$, and aggregates the resulting scores across dialogues. Given a user-proxy agent $\theta_{u}$, the output $R_{\theta_{u}}$ is a structured evaluation record containing per-sample metric values, aggregated results with confidence intervals, and auxiliary metadata such as random seeds, model identifiers, and telemetry reports. We also report metric-level calibration and reliability diagnostics that quantify metric stability and judge variability.
3.2. Datasets & Tasks
We evaluate user proxies across four open-source conversational datasets spanning diverse domains and interaction patterns, collectively providing 795 conversations with real human user turns, enabling direct comparison between proxy-generated and real user behavior.
Table 1. Datasets used for the evaluation. All datasets are open-source and include real human user utterances.
Dataset Samples Avg Turns
ChatbotArena (Zheng et al., 2023) 195 2.5
ClariQ (Aliannejadi et al., 2021) 200 7.0
OASST1 (Köpf et al., 2023) 200 3.3
QULAC (Aliannejadi et al., 2019) 200 2.0
In Table 1, notice that the number of samples we use for benchmarking does not exhaust all the samples of the underlying open-source datasets. Instead, we perform stratified sampling over the original datasets to extract the final evaluation datasets.
For each dataset, we first normalize conversations into alternating user-assistant turn sequences, retaining only English dialogues with at least two turns. We then define dataset-specific stratification keys to ensure broad coverage: OASST1 & ChatbotArena conversations are bucketed by the number of user turns (short, medium, long), QULAC entries are grouped by topic type & facet category, and ClariQ dialogues are distributed across topic buckets & clarification-pair counts. Within each stratum, we allocate samples proportionally to the population size while enforcing a minimum-per-stratum threshold to prevent underrepresentation. Using a fixed random seed for reproducibility, we sample up to 200 samples per dataset, yielding a balanced benchmark that spans diverse conversational structures, topic domains, and interaction patterns without oversampling any single mode.
For each conversation in all four datasets, we generate a user goal description using an auxiliary LLM ($\theta_{g}$ defined in § 3.1) that summarizes the user’s intent, behavior, tone, and persona based on the real conversation. This description serves as the initialization prompt for user proxies during evaluation.
3.3. Metrics
MirrorBench quantifies the human-likeness of user proxies with two metric families: (i) lexical-diversity metrics measure vocabulary richness and repetition patterns (§ 3.3.1), and (ii) judge-based realism metrics measure behavioral realism (Section 3.3.2). All metrics are human-anchored: proxy scores are compared against the empirical distribution of real-user utterances from the same dataset and under same tokenization, yielding interpretable, calibrated results (implementation details in Appendix A.3).
3.3.1. Lexical Diversity Metrics
Lexical-diversity metrics capture how varied a user proxy’s word choices are and how often they repeat tokens. Because raw values are sensitive to sequence length, domain, and tokenization, we normalize ($z$-score) proxy scores with respect to the human distribution on the same dataset.
Moving Average Type-Token Ratio (MATTR)
The classical type–token ratio (TTR) (Chotlos, 1944; Heaps, 1978) decreases as sequence length grows. The moving-average TTR (MATTR) mitigates this length bias by averaging TTR over a sliding window of width $w$ (Covington and McFall, 2010). Let $\text{types}(\cdot)$ return the set of distinct token types in its argument. For a token sequence $\mathbf{t}=(t_{1},\ldots,t_{N})$ and window size $w\leq N$,
(1) $\mathrm{MATTR}_{w}(\mathbf{t})\;=\;\sum_{i=1}^{N-w+1}\frac{\bigl\lvert\mathrm{types}(t_{i},\ldots,t_{i+w-1})\bigr\rvert}{w(N-w+1)},$
where $\mathrm{types}(\cdot)$ returns the set of distinct token types in the window. We use $w=50$ by default unless stated otherwise.
Yule’s K
Yule’s characteristic constant $K$ (Tanaka-Ishii and Aihara, 2015) summarizes repetitiveness from the token-frequency spectrum. Let $V_{i}$ be the number of types occurring exactly $i$ times, and $N=\sum_{i\geq 1}iV_{i}$ the token count. Then
[…]
Lower $K$ indicates richer, less repetitive text; higher $K$ indicates greater repetition. The factor $10^{4}$ provides a convenient numeric range.
Hypergeometric Distribution Diversity (HD-D)
HD-D (Carmona et al., 2019) estimates vocabulary diversity via hypergeometric sampling, offering robustness to length variation. It approximates the expected number of distinct types observed in a sample of size $s$ drawn without replacement from a token sequence $\mathbf{t}=(t_{1},\ldots,t_{N})$. For a type $i$ with frequency $f_{i}$ in $\mathbf{t}$, the probability that the sample contains at least one instance of that type is
[…]
and the HD-D score is their normalized sum:
(3) $\mathrm{HD\text{-}D}_{s}(\mathbf{t})\;=\;\frac{1}{s}\sum_{i=1}^{V}\left(1\;-\;\frac{\binom{N-f_{i}}{\,s\,}}{\binom{N}{\,s\,}}\right)$,
where $V=\bigl|\text{types}(\mathbf{t})\bigr|$. Following the VOCD-D convention, we use $s=42$ by default (Mccarthy and Jarvis, 2007).
Human-Anchored Z-Score Normalization
Raw lexical-diversity scores are dataset-dependent; we therefore report human-anchored $z$-scores computed on the same split and tokenizer. Let $\mathcal{H}=\{\mathbf{t}_{1}^{\mathrm{ref}},\ldots,\mathbf{t}_{n}^{\mathrm{ref}}\}$ be human user-side token sequences, where each $\mathbf{t}_{i}^{\mathrm{ref}}$ concatenates the human user turns of one reference dialogue. For any $m\in\{\mathrm{MATTR}_{w},K,\mathrm{HD\text{-}D}_{s}\}$, let $\mu_{m}^{\text{hum}}=\frac{1}{|\mathcal{H}|}\sum_{\mathbf{t}\in\mathcal{H}}m(\mathbf{t})$ and $\sigma_{m}^{\text{hum}}=\sqrt{\frac{1}{|\mathcal{H}|-1}\sum_{\mathbf{t}\in\mathcal{H}}\bigl(m(\mathbf{t})-\mu_{m}^{\text{hum}}\bigr)^{2}}.$ During evaluation, each rollout episode produces a synthetic proxy-assistant dialogue, which we denote by $\hat{d}_{i}$ (§ 3.1). From $\hat{d}_{i}$, we extract only the proxy user turns (i.e., the outputs of the user–proxy agent $\theta_{u}$) and concatenate them in temporal order to obtain a proxy user token sequence $\hat{\mathbf{t}}_{i}$. This $\hat{\mathbf{t}}_{i}$ is the proxy-side analogue of the human sequence $\mathbf{t}_{i}^{\text{ref}}$ in $\mathcal{H}$. We then assign that rollout episode a human-anchored standardized lexical score:
$z_{m}(\hat{\mathbf{t}}_{i})\;=\;\frac{m(\hat{\mathbf{t}}_{i})-\mu_{m}^{\text{hum}}}{\sigma_{m}^{\text{hum}}}.$
By design, $z_{m}\approx 0$ indicates that the proxy’s lexical behavior matches the human mean for metric $m$. The sign of $z_{m}$ is metric-dependent: for $\mathrm{MATTR}_{w}$ and $\mathrm{HD\text{-}D}_{s}$, $z_{m}>0$ corresponds to greater lexical diversity than the human average; for Yule’s $K$, $z_{m}>0$ corresponds to more repetition (i.e., lower diversity). For reporting at the dataset level, we aggregate across all rollout dialogues $\{\,\hat{d}_{i}\in\hat{D}\,\}_{i=1}^{n}$ executed under a fixed configuration. Let $\{z_{m}(\hat{\mathbf{t}}_{1}),\ldots,z_{m}(\hat{\mathbf{t}}_{n})\}$ be the per-episode standardized scores for metric $m$. We report their sample mean $\bar{z}_{m}$ together with a two-sided $95\%$ confidence interval derived from the Student-$t$ distribution:
(4) $\mathrm{CI}_{0.95}(m)\;=\;\bar{z}_{m}\;\pm\;t_{0.975,\,n-1}\,\frac{s_{m}}{\sqrt{n}},$
where $s_{m}$ is the unbiased sample standard deviation, and $t_{0.975,\,n-1}$ is the $97.5$th percentile of the $t$-distribution with n1n-1 degrees of freedom. These aggregated statistics, namely $\bar{z}_{m}$, $s_{m}$, and $\mathrm{CI}_{0.95}(m)$, are the quantities that MirrorBench reports for downstream comparison across user proxies, datasets, and seeds.
3.3.2. Judge-Based Realism Metrics
Lexical diversity alone cannot capture whether a simulated user feels human-like. Human-likeness also depends on discourse phenomena such as tone, politeness, hesitations, or style. MirrorBench therefore includes judge-based realism metrics: LLM evaluators that score proxy behavior against human references along higher-level dimensions. Each judge call uses a chain-of-thought style prompting strategy (Wei et al., 2022): the judge is instructed to reason privately and then produce a final structured verdict. To improve robustness, each judge is executed with a self-consistency parameter $c\geq 1$, which repeats the judgment $c$ times under the same prompt and aggregates the resulting scores.
GTEval (Relative Realism Scoring)
GTEval (Zhu et al., 2025) compares a proxy-generated user transcript against a human reference and returns a similarity score in $[0,1]$, where higher values indicate that the proxy is judged to behave more like the human reference. The judge model (with $c=1$) is shown the pair $(\hat{d}_{j},d_{j}^{\mathrm{ref}})$ and asked to assess stylistic and behavioral similarity (e.g., tone, naturalness). Let $\mathrm{J}^{(r)}(\cdot)\in[0,1]$ denote the scalar verdict returned by the judge on repetition $r$. The GTEval score for episode $j$ is
(5) $s^{(\mathrm{GTEval})}_{j}\;=\;\frac{1}{c}\sum_{r=1}^{c}\mathrm{J}^{(r)}\!(\hat{d}_{j},\,d_{j}^{\mathrm{ref}}),$
where each $\mathrm{J}^{(r)}$ corresponds to an independent call to the judge with a different random seed.
Pairwise Indistinguishability (PI)
Here, judge LLM is given two anonymized user-assistant conversations, and asked to choose one in which the user sounds more like a human. For each sample, we present one proxy conversation $\hat{d}_{j}$ and one human reference $d_{j}^{\mathrm{ref}}$ in random order, unlabeled except as $\mathrm{A}$ or $\mathrm{B}$. The judge returns a verdict $v_{j,r}\in\{\mathrm{A},\mathrm{B},\mathrm{Tie}\}$ for judgment round $r$. Let $p_{j,r}\in\{\mathrm{A},\mathrm{B}\}$ be the position where the proxy was shown for that comparison (randomized per judgment). With $c$ repeated judgments (default $c=3$), we define the per-episode win rate
(6) $w^{(\mathrm{PI})}_{j}\;=\;\frac{1}{c}\sum_{r=1}^{c}(\mathds{1}[v_{j,r}=p_{j,r}]\;+\;0.5\,\mathds{1}[v_{j,r}=\mathrm{Tie}]).$
Aggregated win rate is summarized by the mean $\bar{w}^{(\mathrm{PI})}=\frac{1}{n}\sum_{j=1}^{n}w^{(\mathrm{PI})}_{j}$ and its centered form $\Delta w^{(\mathrm{PI})}=\bar{w}^{(\mathrm{PI})}-0.5$. Here, $\Delta w>0$ indicates the judge tends to prefer the proxy user over the human baseline, $\Delta w<0$ favors real users, and $\Delta w\approx 0$ indicates that the proxy is effectively indistinguishable from the human baseline.
Rubric-and-Reason (RNR)
RNR adapts rubric-style LLM evaluation to judge human-likeness without requiring a paired human reference. The judge sees only the proxy-side user transcript $\hat{d}_{j}$ and a rubric that defines realism along dimensions such as style, behavior, and tone. For a given rubric, the judge returns a binary verdict $v^{(r)}\in\{\mathrm{YES},\mathrm{NO}\}$ on repetition $r$, which we map to
$\hat{s}_{\mathrm{RNR}}^{(r)}(\hat{d}_{j})\;=\;\begin{cases}1.0,&\text{if }v^{(r)}=\mathrm{YES},\\ 0.0,&\text{if }v^{(r)}=\mathrm{NO}.\end{cases}$
We repeat each judgment $c=2$ times. The aggregate RNR score for sample $j$ is
(7) $s_{\mathrm{RNR}}(\hat{d}_{j})\;=\;\frac{1}{c}\sum_{r=1}^{c}\hat{s}_{\mathrm{RNR}}^{(r)}(\hat{d}_{j}).$
Unlike GTEval and PI, which are explicitly comparative, RNR is reference-free and produces an absolute realism score for the proxy alone. Because this absolute score depends on the judge model and the elaborated rubric, it is interpreted relatively: higher $s_{\mathrm{RNR}}(\hat{d}_{j})$ means the proxy is behaving more like a real user under the fixed judge and rubric set. Optionally, we also evaluate the human reference $d_{j}^{\mathrm{ref}}$ under the same procedure to obtain $s_{\mathrm{RNR}}(d_{j}^{\mathrm{ref}})$, which serves as an empirical upper bound (useful for calibration).
Figure 2. Human-likeness of five user-proxy LLMs across four datasets. Higher is better for judge-based metrics (GTEval, PI, RNR). Lexical-diversity metrics are $z$-scored to human baselines (0 is best). We fix the judge to Claude-4-Sonnet and the assistant to GPT-4o.
Calibration via HH and PP Controls
Judge models can exhibit systematic bias (e.g., favoring verbosity, excessive politeness, or self-preference). To expose this bias and to place GTEval and PI scores on an interpretable scale, we optionally compute two control conditions:
  • Human-Human (HH): Compare a human conversation to itself, $s_{\mathrm{HH},j}=\mathrm{Judge}(d_{j}^{\mathrm{ref}},d_{j}^{\mathrm{ref}})$. This approximates the judge’s effective ceiling for genuine human behavior on episode $j$.
  • Proxy-Proxy (PP): Compare a proxy conversation to itself, $s_{\mathrm{PP},j}=\mathrm{Judge}(\hat{d}_{j},\hat{d}_{j})$. This estimates how the judge scores the proxy in isolation.
For GTEval, both $s_{\mathrm{HH},j}$ and $s_{\mathrm{PP},j}$ are typically expected to be close to 1. For PI, a self comparison should yield an ideal effective win rate near 0.5. We calculate their empirical means across all $n$ samples: $\mu_{\mathrm{HH}}=\tfrac{1}{n}\sum_{j=1}^{n}s_{\mathrm{HH},j},$ $\mu_{\mathrm{PP}}=\tfrac{1}{n}\sum_{j=1}^{n}s_{\mathrm{PP},j}.$ For PI Metric, we also calculate calibrated score $s_{\mathrm{cal}}$. Let $s_{\mathrm{raw}}$ denote the uncalibrated proxy score of interest (e.g. $\bar{w}^{(\mathrm{PI})}$). We define a calibrated $[0,1]$ score via an affine rescaling between the PP and HH anchors:
(8) $s^{PI}_{\mathrm{cal}}\;=\;\mathrm{clip}\!\left(\frac{s_{\mathrm{raw}}-\mu_{\mathrm{PP}}}{\max\!\bigl(\epsilon,\;\mu_{\mathrm{HH}}-\mu_{\mathrm{PP}}\bigr)},\;0,\;1\right),$
where $\epsilon=10^{-6}$ prevents division by zero and $\mathrm{clip}(\cdot,0,1)$ bounds the result. Under this normalization, $s_{\mathrm{cal}}\approx 0$ means the proxy is no better than its own PP baseline, while $s_{\mathrm{cal}}\approx 1$ means the proxy is judged as indistinguishable from human under the same judge.
4. Experiments & Results
We evaluate multiple user-proxy LLMs across four diverse conversational datasets, ChatbotArena, ClariQ, OASST1, and QULAC, to quantify human-likeness using both lexical-diversity and LLM-judge realism metrics. Concretely, we report MATTR, Yule’s $K$, and HD-D (human-anchored $z$-scores), together with three judge-metrics GTEval, PI, and RNR. Our experiments (i) measure how closely different proxy architectures reproduce human user behavior, (ii) assess the reliability and calibration of LLM-as-judge metrics, and (iii) characterize computational cost, latency, and throughput under large-scale runs.
We compare five LLMs as user proxies: GPT-4o (Hurst et al., 2024), GPT-5 (OpenAI, 2025a), GPT-OSS-120B (OpenAI, 2025b), Claude-4-Sonnet (Anthropic, 2025), and Gemini-2.5-Pro (Google, 2025). Unless otherwise noted, the assistant is fixed to GPT-4o and the judge to Claude-4-Sonnet for all primary results, enabling apples-to-apples comparisons across proxies. All the evaluations are performed with a single seed (unless otherwise specified), and we report aggregated scores with 95% confidence intervals. Judge scores are calibrated, and we include a human correlation check to validate judge trends. We also analyze fine-grained telemetry to make cost/performance trade-offs explicit.
4.1. Comparison of User-Proxy LLMs
Figure 2 compares five user-proxy LLMs across four datasets with a fixed judge (Claude-4-Sonnet) and assistant (GPT-4o). The left three columns of subplots report judge-based realism (GTEval, PI $\Delta w$, RNR; higher is better); the right three columns report human-anchored lexical diversity (MATTR, HD-D, Yule’s K; best at the human z-score = 0). Dashed red (HH) and green (PP) lines provide calibration controls; error bars denote 95% CIs. We provide the assistant the reference chat history to anchor the rollout to the same trajectory as the original conversation; while LLM hallucinations can occasionally induce drift, a manual inspection found such divergence in mere $1\%$ of samples evaluated as part of 2.
Judge realism is consistent across datasets
Across GTEval, PI $\Delta w$, and RNR, Gemini-2.5-Pro and Claude-4-Sonnet are the most human-like on every dataset, with GPT-4o competitive but generally behind, and GPT-OSS-120B and GPT-5 trailing. On ClariQ and QULAC, Claude-4-Sonnet/Gemini-2.5-Pro approach the HH ceiling under RNR, while PI $\Delta w$ shows clear positive win margins for these models and negative/near-zero deltas for the rest. The agreement among the three judge metrics indicates a stable ordering not driven by any single rubric.
Diversity shows strong dataset effects and a realism–diversity tension
Lexical diversity diverges from the realism ranking. On ClariQ, most notably Claude-4-Sonnet and GPT-5 exceed the human anchor on MATTR/HD-D and exhibit lower Yule’s K, indicating a more diverse vocabulary than human questioners in this information-seeking regime. In contrast, QULAC exhibits a uniform diversity deficit: all proxies fall below the human baseline on MATTR and HD-D (with positive shifts in Yule’s K), suggesting more templated clarifications than humans. ChatbotArena and OASST1 sit between these extremes with smaller deviations around the human baseline. Overall, Gemini-2.5-Pro & GPT-4o yield the strongest diversity alignment with humans across datasets, whereas Claude-4-Sonnet generally tends to overshoot the human baseline. GPT-5 and GPT-OSS-120B have generally larger diversity difference compared to human baselines.
Judge realism and diversity are partially decoupled
High judge realism does not guarantee human-level diversity. Claude-4-Sonnet and Gemini-2.5-Pro lead on GTEval/PI/RNR but under-shoot diversity on QULAC, indicating judges favor intent/style over surface variety. Conversely, GPT-4o shows more stable diversity with moderate judge-realism gains, underscoring that diversity alone is insufficient to achieve judge indistinguishability.
Uncertainty and robustness.
Confidence intervals are generally narrow. When overlaps occur (e.g., GTEval on ChatbotArena), rank differences are small and concur with PI/RNR trends. For lexical-diversity panels, confidence intervals occasionally widen, notably for Yule’s K, due to very short the user-side text (e.g., due to few turns) that increases estimator variance. Overall, Figure 2 supports three takeaways: (i) Gemini-2.5-Pro and Claude-4-Sonnet are reliably the most human-like by judge criteria; (ii) diversity depends strongly on dataset regime and proxies often lag on clarification-centric tasks like QULAC; and (iii) realism and diversity capture complementary facets of user-proxy quality, motivating the use of both families of metrics.
4.2. Judge Sensitivity Analysis
With the assistant and user-proxy fixed to GPT-4o, Figure 3 shows substantial judge sensitivity on realism scores. On GTEval, scores spread widely ($\approx\!0.45\text{--}0.81$), with GPT-4o as judge yielding the highest value. PI is the most volatile: Gemini-2.5-Pro & Claude-4-Sonnet produce near-zero or negative win deltas, while GPT-5 & GPT-4o are clearly positive, suggesting family/self-preference or rubric alignment effects. RNR saturates near the ceiling for GPT-4o & GPT-5 judges ($\approx\!0.96\text{--}0.98$) and is lower for others ($\approx\!0.79\text{--}0.82$), indicating judge-dependent sensitivity: PI >> GTEval >> RNR.
Figure 3. Judge sensitivity of judge-realism metrics on ChatbotArena with assistant & user-proxy both set to GPT-4o; bars vary the judge model. Error bars are 95% CIs.
These differences imply that conclusions drawn from a single judge can shift both the absolute level and the ordering of models. Since LLM judges can exhibit systematic biases (Thakur et al., 2025; Sahoo et al., 2025), they may behave either overly generous or overly conservative. Thus in practice, we recommend evaluating with multiple judges and applying HH/PP calibration when feasible, or normalizing per-judge before aggregation, to avoid over-interpreting judge-specific biases. Empirically, Claude-4-Sonnet behaves as a conservative yet stable judge, yielding non-saturated, well-separated scores across metrics, which is why we use Claude-4-Sonnet as a judge in the results reported in Figure 2.
4.3. Human-Judge Correlation
Although judge-sensitivity analysis (§4.2) helps characterize a judge’s behavior on our task, it can still be difficult to interpret, since whether a judge appears “generous” or “conservative” is itself task-dependent. We therefore additionally assess judge reliability via human-judge correlation.
Figure 4. Judge–human correlation on ChatbotArena: Correlation of Claude-4-Sonnet judge scores & human scores for GTEval and PI, evaluated on Gemini-2.5-Pro user-proxy outputs (N=100 per metric). Solid line: linear fit; dashed: identity. Point size encodes local sample density. All correlations p << 0.001.
We validate judge reliability by correlating Claude-4-Sonnet judge scores with blinded human expert annotations on ChatbotArena. We stratify 100 episodes per metric (GTEval, PI) by judge score and conversation length, then compute Spearman’s $\rho$ (Kokoska and Zwillinger, 2000), Pearson’s $r$ (Kowalski, 2018), and Kendall’s $\tau$ (Kendall, 1938). As shown in Figure 4, GTEval aligns strongly with human judgments, while PI exhibits a moderate correlation. Note that PI is inherently harder to annotate because it requires pairwise comparisons; thus, a moderate correlation still indicates that PI is a reasonably reliable judge metric. Overall, these results suggest that judge scores track human perceptions of user-proxy quality.
4.4. Assistant Sensitivity and Rank Stability
LLM-as-judge scores can be confounded by the assistant model that the user proxy interacts with. To quantify this assistant sensitivity, we rerun evaluation on ChatbotArena while holding the judge fixed and varying the assistant among {GPT-4o, Claude-4-Sonnet, Gemini-2.5-Pro}. For each user-proxy model, we compute the three judge-based realism metrics (GTEval, PI, RNR) on the resulting rollouts.
Figure 5. Assistant sensitivity on ChatbotArena (scores). Mean $\pm$ 95% CI of judge-based realism metrics for each user-proxy when swapping the assistant model (color); the judge is fixed to Claude-4-Sonnet.
Figure 6. Rank stability under assistant swaps (ChatbotArena). Per-metric proxy ranks (1 = best) induced by the scores in Fig. 5 when varying the assistant model; stable lines indicate robust proxy ordering, while crossings reveal assistant-dependent rankings.
Figure 5 shows that absolute scores vary only slightly across assistants and are generally insufficient to change the overall ordering of user-proxy models. Figure 6 makes this explicit: ranks are stable (largely parallel trajectories) across assistants, with a single crossover in GTEval. Overall, assistant choice introduces mild variance in judge-based realism scores, but the resulting rankings remain stable most of the time.
4.5. Multi-Seed Robustness
LLM-as-judge metrics can fluctuate across repeated runs due to stochasticity in generation and evaluation. To assess the stability of our judge-based realism metrics, we perform a multi-seed analysis by repeating the benchmark five times with different random seeds for three different user-proxies while keeping the assistant and judge fixed.
Figure 7. Multi-seed robustness of judge-based realism metrics. For each user-proxy model, points show the per-seed mean score and error bars indicate the corresponding 95% CI on ChatbotArena; the assistant (gpt-4o) and judge (claude-4-sonnet) are fixed across runs.
[…]
4.6. Telemetry, Scaling, and Cost
Based on our experiments, we further provide detailed analysis on the practical side of evaluation. Since most of the models we used are proprietary (incurs cost per million tokens), we provide statistical analysis on cost, latency and trade-off between cost vs performance. Per-sample telemetry (Fig. 8) shows that token usage is dominated by the judge, with the user-proxy & assistant contributing a smaller but non-negligible share; datasets differ markedly, OASST1 drives the highest token counts while ClariQ yields the largest end-to-end latency.
Figure 8. Avg per-sample telemetry for GPT-4o as user-proxy & assistant, and judged by Claude-4-Sonnet across 4 datasets on 6 metrics (Fig. 2). Left: Token usage by role (darker: input, lighter: output). Right: Cumulative latency per episode.
Figure 9. Cost-quality trade-off for PI: cost per evaluation (USD) vs. PI $\Delta w$ ($\uparrow$ better). Judge = Claude-4-Sonnet; assistant = GPT-4o; temperature = 0; cache off. Markers denote user-proxies; labels denote datasets. Dashed line: Pareto frontier. See Table 4 for model pricing.
The cost-quality frontier (Fig. 9) clarifies trade-offs for PI evaluation: User-proxies Gemini-2.5-Pro and Claude-4-Sonnet offer attractive Pareto points (good PI $\Delta w$ at moderate cost), and GPT-5 generally incurs higher cost with weaker PI gain. Overall, user-proxy Gemini-2.5-Pro, along with Claude-4-Sonnet as judge, provides a balanced choice for large runs.
5. Conclusion & Future Work
MirrorBench provides a reproducible benchmark for evaluating the human-likeness of LLM-based user proxies, assessing how closely a proxy’s utterances match real human users across conversational tasks, independent of downstream task success. Across four public datasets, we evaluate five user-proxy LLMs using three judge-based realism metrics (GTEval, PI, RNR) together with human-anchored lexical diversity measures (MATTR, HD-D, Yule’s K). The resulting analyses highlight (i) systematic gaps between proxies and human users, (ii) a consistent realism-diversity trade-off that varies by dataset regime, and (iii) measurable sensitivity of absolute judge scores to the judge model, motivating calibration controls (HH/PP) and transparent multi-judge reporting. We also report token, latency, and cost profiles to make the evaluation trade-offs explicit when benchmarking at scale.
Limitations and outlook. Judge-based metrics may reflect model-family bias and prompt sensitivity; while HH/PP controls provide useful context, they do not eliminate all sources of bias. Our experiments primarily use a fixed assistant configuration and limited randomization, and our coverage is restricted to four English-centric datasets. Finally, lexical diversity captures surface-level distributional variation and does not fully represent interaction-level diversity such as repair, initiative, or long-horizon consistency.
Overall, MirrorBench is intended as a benchmark for systematically comparing user proxies and for grounding progress on realistic user simulation in conversational AI.
Appendix A MirrorBench System
In this section, we explain the system aspects of the benchmarking framework. Although the metric definitions and execution may seem straightforward, substantial work went into creating a scalable, extensible system for flexible research use. We provide a detailed architectural blueprint in §A.1, explain the execution flow in §A.2, and provide metric implementation details in §A.3
A.1. Architecture
Figure 10. MirrorBench Architecture: Six-layer stack from low-level execution backends & persistence up through the core engine, plugin components, CLI & reporting, and task drivers. Top layers are user-facing; bottom layers are low-level infrastructure abstractions.
MirrorBench is a six-layer stack that cleanly separates infrastructure from evaluation logic (Fig. 10). Layers build upward, lower layers provide execution and data services; upper layers host pluggable components, task logic, and user interfaces.
A.1.1. Execution Backends & Persistence
This layer is responsible for all the execution and data management. The layer consists of two high-level modules explained below.
Execution Backend
It is responsible for taking decomposed units of tasks as input and send them for execution on the respective execution engine. The framework provides abstractions to support multiple execution backends to execute evaluation jobs at scale. The abstract interface provides uniform schema for implementations to extend the support for new execution engines. The framework supports fully-functional synchronous & asynchronous execution backends, and provides stub implementation to extend for Ray (Moritz et al., 2018) backend.
Persistence
All evaluation artifacts are persisted in a SQLite database organized around a three-level hierarchy: runs define experiment configurations, units specify individual evaluation tasks (parameterized by user-proxy, dataset, metric, and seed), and episodes capture per-trial results including execution duration, metric values, and telemetry. Aggregate statistics (mean, standard deviation, confidence intervals) are materialized in a separate metrics table to support efficient querying, while composite scores across multiple metrics are stored in scorecards for holistic comparisons. The complete schema is detailed in Table 2.
Table 2. Complete MirrorBench Run Database Schema
Table Field Type Description
runs run_id TEXT Unique identifier for the experiment run (primary key)
created_at TEXT ISO-8601 timestamp of run creation
status TEXT Run status: created, running, completed, failed
engine TEXT Execution engine identifier (e.g., sync, async)
planner_version TEXT Version of the planner used for the run
summary_json TEXT JSON-encoded run-level summary statistics
notes TEXT User-provided annotations or metadata
units run_id TEXT Foreign key to parent run
unit_id TEXT Unique identifier for the evaluation unit (composite primary key)
user_proxy TEXT User-proxy agent implementation identifier
dataset TEXT Dataset name (e.g., oasst1, clariq)
metric TEXT Evaluation metric (e.g., gteval)
seed INTEGER Random seed for reproducibility
judge TEXT Optional judge model for LLM-based metrics
status TEXT Unit execution status: pending, running, completed
episodes run_id TEXT Foreign key to parent run
unit_id TEXT Foreign key to parent unit
episode_id TEXT Unique episode identifier (primary key)
status TEXT Episode outcome: success, failure, timeout
duration_s REAL Wall-clock execution time in seconds
artifact_path TEXT File path to serialized conversation artifact
summary TEXT Human-readable episode summary
metric_values TEXT JSON-encoded per-metric scores for this episode
telemetry_json TEXT JSON-encoded execution telemetry (token counts, API calls)
metrics run_id TEXT Foreign key to parent run
unit_id TEXT Foreign key to parent unit
metric TEXT Registered metric name (primary key)
mean REAL Sample mean across episodes
standard_deviation REAL Sample standard deviation
confidence_interval REAL 95% confidence interval half-width
p_value REAL Statistical significance (e.g., vs. baseline)
sample_size INTEGER Number of valid episodes
extras TEXT JSON-encoded additional statistics
telemetry run_id TEXT Foreign key to parent run
unit_id TEXT Foreign key to parent unit
key TEXT Telemetry key (e.g., total_tokens, api_errors)
value TEXT Aggregated telemetry value (JSON or scalar)
scorecards run_id TEXT Foreign key to parent run
name TEXT Scorecard identifier (e.g., realism, diversity)
score REAL Weighted composite score
weights TEXT JSON-encoded metric weights used for aggregation
missing_metrics TEXT JSON list of metrics excluded due to missing data
extras TEXT JSON-encoded additional scorecard metadata
A.1.2. Core Engine
All the core functionalities like data models, registry builder, configuration and manifest management are contained in Layer 2 of the architecture. Since Layer 1 provides uniform abstractions to produce and consume objects, this layer considers backend-agnostic implementation.
Data Models
All the data structures for messages, episodes, evaluation units and run manifests are defined with typed data models using TypedDict and pydantic. We formally explain the context of these different data objects later.
Registry Builder
The framework provides certain high-level components which requires module-specific registries. This layer provides the abstract registry builder that supports creating metadata-aware registry for the downstream modules. Such registries can avoid silent errors since we can ensure certain constraints on the classes being registered with the metadata.
Configuration & Manifest Management
A typical evaluation job requires configuring user-proxy, datasets, metrics, scorecards etc. We define standard configuration schema for each of such components. Moreover, upon invocation of job or during dryrun, the system decomposes evaluation task into small evaluation units. In order to support parallelization (for async and remote backends), we persist self-contained information about the executable units as manifest files.
A.1.3. Orchestration
This layer coordinates the evaluation workflow by managing component lifecycles, execution planning, and runtime operations. The layer consists of five orchestration modules.
Pipeline Planner
It validates job configurations and generates execution plans by decomposing evaluation jobs into granular evaluation units (user-proxy, dataset, metric, seed combinations). The planner performs compatibility filtering based on component metadata to ensure valid combinations and produces reproducible plan manifests.
Component Registries
It maintains metadata-aware registries for all pluggable components including user-proxy adapters, datasets, metrics, model clients, and judges. Registries support dynamic component discovery and instantiation through decorator-based registration patterns.
Run Controller
It orchestrates end-to-end evaluation runs by managing execution flow, coordinating backend selection (synchronous, asynchronous, or Ray), tracking progress, and handling run lifecycle including resumption of interrupted runs.
Cache Layer
It provides SQLite-based caching for model responses to reduce redundant API calls and improve evaluation efficiency. The cache uses content-based keys (hashing model name, messages, and configuration) and supports namespace isolation and TTL-based expiration.
Observability Layer
It provides structured logging via structlog and optional OpenTelemetry integration for distributed tracing and metrics collection. The layer supports multiple output formats (JSON, text) and configurable log levels for debugging and monitoring.
A.1.4. Plugin Components
This layer contains the core pluggable components that implement domain-specific evaluation logic. All components follow standardized interfaces and register through the registry system.
Model Clients
It wraps provider SDKs (OpenAI, LangChain, Azure, etc.) with a unified interface for LLM invocation. Clients automatically capture telemetry (tokens, latency, cost) and support transparent caching through wrapper composition.
User-Proxy Adapters
It adapts diverse agent frameworks and LLM configurations into a standardized AgentSession interface. Adapters normalize differences in message formats, tool usage patterns, and execution models across frameworks.
Datasets
It loads and normalizes evaluation datasets from various sources (HuggingFace, JSONL files, etc.) into standardized episode objects. Loaders support dataset splitting, sampling limits, and provide metadata about available splits and reference statistics.
Metrics
It implements human-likeness measurements including lexical diversity metrics (MATTR, HD-D, Yule’s K) and LLM-as-judge metrics (GTEval, PI, RNR). Metrics declare compatibility requirements through metadata (task types, judge dependencies, reference statistics needs).
A.1.5. Task Drivers
This layer implements interaction protocols that orchestrate how the artificial conversation between user-proxy & assistant agent is synthesized and how it is engaged with the evaluation metrics. Task drivers bridge the gap between high-level evaluation specifications and low-level execution.
Default Single-Turn Task Driver
It executes single-turn interactions where the user-proxy receives episode context and generates a single response. This driver is suitable for simple prompt-completion tasks and turn-level utterance evaluation.
Mirror Conversation Task Driver
It orchestrates multi-turn conversational interactions by executing sequences of user-proxy responses within episodic contexts. This driver supports conversation-level evaluations where human-likeness is assessed across extended interactions.
A.1.6. API & Interface
This layer provides user-facing interfaces for interacting with the framework through programmatic APIs and command-line tools.
Runner Facade
It provides a simplified programmatic API for executing evaluation runs from Python code. The facade abstracts orchestration complexity and exposes high-level methods for planning, execution, and result retrieval with progress tracking.
Command-Line Interface (CLI)
It exposes framework functionality through a comprehensive CLI supporting workflow commands (plan, dryrun, run), reporting (report json), run management (runs inspect, runs delete), and cache operations (cache stats, cache purge).
Report Generation
It produces structured evaluation summaries with aggregated statistics including means, confidence intervals, and significance testing. The module supports JSON report output with planned support for HTML visualization.
A.2. Execution Flow
Figure 11. MirrorBench execution flow. The framework decomposes an evaluation job into units $\{U_{1},\ldots,U_{p}\}$; each unit $U_{i}$ iterates over episodes $\{e_{1},\ldots,e_{n}\}$ produced by the dataset adapter and executed via task drivers. Metrics are computed per episode and aggregated within each unit with confidence intervals.
This section details the end-to-end execution that instantiates the pipeline $\Psi_{T}$ (as described in §3.1). Figure 11 shows the path from CLI invocation through planning, orchestration, and nested execution to persisted results, compiling high-level specifications into reproducible evaluation artifacts.
CLI Invocation and Planning
An evaluation job is declared via a configuration file (appendix §B.2) that specifies user proxies $\mathcal{P}=\{\theta_{u_{1}},\ldots,\theta_{u_{q}}\}$, datasets $\mathcal{D}=\{D_{1},\ldots,D_{d}\}$, the metric set $M=\{m_{1},\ldots,m_{k}\}$, a fixed assistant model $\theta_{a}$, and run parameters (backend, observability, maximum concurrency, and a seed set $S=\{s_{1},\ldots,s_{r}\}$). The seed set enables repeated trials of the same combination under distinct random initializations, supporting variance estimation and confidence intervals.
The Planner validates the configuration against component registries and enumerates all mutually compatible tuples to produce execution units $U=\{U_{1},\ldots,U_{p}\}$.
$U=\left\{(\theta_{u_{x}},D_{y},m_{z},s_{w})\,\middle|\,\begin{aligned} &\theta_{u_{x}}\in\mathcal{P},\;D_{y}\in\mathcal{D},\\ &m_{z}\in M,\;s_{w}\in S,\\ &\operatorname{Compat}(\theta_{u_{x}},D_{y},m_{z})\end{aligned}\right\}$
Here $\operatorname{Compat}(\cdot)$ returns true iff the components can be jointly executed under their declared capabilities and requirements (e.g. metric prerequisites, model/back-end constraints). Each unit $U_{i}=(\theta_{u_{x}},D_{y},m_{z},s_{w})$ is executed and evaluated independently, with models, prompts, and driver assignments fixed at plan time. The planner creates & persists a replayable manifest file (appendix §B.2) capturing all requisite metadata, and exposes a dryrun checkpoint for inspection prior to committing computational resources.
Orchestration and Backend Selection
The Runner instantiates a Run Controller that manages execution state, telemetry, and result persistence. The controller initializes a run-specific SQLite database to store task driver outputs and metric values. Based on the run configuration, the system selects an available backend (sync, async, or distributed) and dispatches the list of units $U$ for execution under the chosen strategy. For each unit $U_{i}\in U$, the backend invokes a unit executor that loads the dataset, instantiates the proxy, initializes the metrics, and configures the task driver.
Unit Execution and Episode Extraction
For each unit $U_{i}=(\theta_{u_{x}},D_{y},m_{z},s_{w})$, the unit executor loads dataset $D_{y}$ and enumerates its constituent episodes. Each episode $e_{j}$ is an immutable specification of a single reference dialogue (§3.1), with dataset layout and fields given by
$D_{y}$ $=\;\{\,e_{1},e_{2},\ldots,e_{n}\,\},$
$e_{j}$ $=\;\bigl(h_{j}^{\mathrm{ref}},\,g_{j},\,\mathrm{meta}_{j}\bigr),$
$h_{j}^{\mathrm{ref}}$ $=\;\bigl[(u_{j,1}^{\mathrm{ref}},a_{j,1}^{\mathrm{ref}}),\ldots,(u_{j,L_{j}}^{\mathrm{ref}},a_{j,L_{j}}^{\mathrm{ref}})\bigr],$
where $g_{j}$ denotes the user goal (explicit or inferred), and $\mathrm{meta}_{j}$ is an optional field containing evaluation-specific annotations (e.g., task tags or reference statistics). The unit executor iterates over the episode set $\{e_{j}\}$, delegating each episode to the appropriate episode executor for execution. Episode executor first synthesizes proxy-assistant conversation using task driver and later compute evaluation results using metric implementation $m_{z}$.
Conversation Synthesis via Task Drivers
For each episode $e_{j}$, the task driver orchestrates the synthesis of a proxy-assistant interaction by alternating invocations of the user proxy $\theta_{u_{x}}$ (abbrev. $\theta_{u}$) and the assistant model $\theta_{a}$, matching the reference dialogue length $L_{j}=|h_{j}^{\mathrm{ref}}|$. Let $\hat{h}_{j,t}$ denote the synthetic history after $t$ turns, with $\hat{h}_{j,0}=\langle\,\rangle$. The synthetic rollout for episode $e_{j}$ is:
$\textbf{Initialize:}\quad\hat{h}_{j,0}\leftarrow\langle\,\rangle$
$\textbf{For }t=1,\ldots,L_{j}\textbf{:}$
$\qquad\hat{u}_{j,t}\;\leftarrow\;\theta_{u}\!\bigl(\hat{h}_{j,t-1},\,g_{j}\bigr),$
$\qquad\hat{a}_{j,t}\;\leftarrow\;\theta_{a}\!\bigl(\hat{h}_{j,t-1}\mathbin{\|}(\hat{u}_{j,t}),h_{j}^{\mathrm{ref}}\bigr),$
$\qquad\hat{h}_{j,t}\;\leftarrow\;\hat{h}_{j,t-1}\mathbin{\|}(\hat{u}_{j,t},\hat{a}_{j,t}).$
The driver records per-turn telemetry (latency and token usage) and produces the episode artifact
$r_{j}\;=\;\bigl(\,\hat{h}_{j,L_{j}},\;e_{j},\;\mathrm{telemetry}_{j}\bigr),$
which binds the synthetic transcript to its reference. The synthetic history is persisted in cache, ensuring that other metric computations can reuse it.
Table 3. Experimental Configuration and Dependency Versions
Category Component Version/Configuration
Hardware & Infrastructure
OS Linux SUSE Linux Enterprise Server 15 SP4 (kernel 5.14+)
CPU Standard Intel Xeon Platinum 8468 (16+ cores)
RAM Standard 64 GB DDR4
[…]
CUDA GPT-OSS-120B only CUDA 12.4+
Runtime & Storage
Python Interpreter 3.12.0+
SQLite Database 3.45.0+ (WAL mode enabled)
API Providers (accessed Oct–Dec 2024)
OpenAI GPT-4o gpt-4o-2024-08-06
GPT-5 gpt-5-2025-08-07
Anthropic Claude-4-Sonnet claude-sonnet-4-20250514
Vertex AI Gemini-2.5-Pro 2025-06-17
Self-hosted GPT-OSS-120B vLLM 0.10.1+gptoss, BF16 precision
Core Library Versions
Model Clients openai 1.40.0+
langchain-openai 0.1.0+
Data pydantic 2.5.0+
datasets 4.1.1+ (HuggingFace)
Utilities tiktoken 0.7.0+ (tokenization)
nltk 3.9.2+ (sentence segmentation)
tenacity 8.2.3+ (retry logic)
Logging structlog 24.1.0+
Telemetry opentelemetry-sdk 1.37.0+
opentelemetry-exporter-otlp 1.37.0+
Metric Computation
After producing an episode artifact $r_{j}$, metric $m_{z}$ evaluates $r_{j}$ to yield a scalar value $v_{j,m_{z}}$. Depending on the metric type (e.g., lexical, or LLM-judge), the computation may involve reference comparison, pairwise scoring, or external model inference. Formally,
$v_{j,m_{z}}\;=\;m_{z}\!\bigl(r_{j}\bigr),$
where $v_{j,m_{z}}\in\mathbb{R}$ represents the realism score for metric $m_{z}$ on episode $e_{j}$. Each evaluation may also produce auxiliary data such as judge verdicts, raw reasoning traces, or calibration diagnostics. The full episode artifact, together with all metric outputs, is persisted to disk for aggregation, post-hoc inspection and reproducibility.
Result Aggregation
After processing all episodes $\{e_{1},\ldots,e_{n}\}$ within unit $U_{i}$, the executor aggregates the metric outputs into summary statistics. For metric $m_{z}$, let $V_{U_{i},m_{z}}=\{v_{1,m_{z}},v_{2,m_{z}},\ldots,v_{n,m_{z}}\}$ denote the set of episode-level scores. The unit-level aggregate includes the mean, standard deviation, and $95\%$ confidence interval:
$\bar{v}_{U_{i},m_{z}}\;$ $=\;\frac{1}{n}\sum_{j=1}^{n}v_{j,m_{z}},$
$s_{U_{i},m_{z}}\;$ $=\;\sqrt{\frac{1}{n}\sum_{j=1}^{n}(v_{j,m_{z}}-\bar{v}_{U_{i},m_{z}})^{2}},$
$\mathrm{CI}_{95}(U_{i},m_{z})$ $=\left[\begin{array}[]{@{}l@{}}\bar{v}_{U_{i},m_{z}}-1.96\,\dfrac{s_{U_{i},m_{z}}}{\sqrt{n}}\\[6.0pt] \bar{v}_{U_{i},m_{z}}+1.96\,\dfrac{s_{U_{i},m_{z}}}{\sqrt{n}}\end{array}\right].$
The run database stores these aggregates, telemetry totals, and per-episode status indicators (success/failure).
Iteration and Reporting
The backend executes each unit $U_{i}=(\theta_{u_{x}},D_{y},m_{z},s_{w})$ in $U=\{U_{1},\ldots,U_{p}\}$. Upon completion, the controller updates the last-run pointer for convenient access and consolidates a unified RunSummary that collates, for every $m_{z}$ and unit $U_{i}$, the tuple $\bigl(\bar{v}_{U_{i},m_{z}},\,s_{U_{i},m_{z}},\,\mathrm{CI}_{95}(U_{i},m_{z})\bigr)$, along with cumulative telemetry and execution metadata. This summary serves as the canonical artifact for subsequent JSON/HTML report generation (appendix §B.2) and cross-proxy comparison. The caching layer eliminates repeated LLM calls across units/runs, and the plan manifest enables exact replays of component instantiation and invocation for reproducible results.
A.3. Implementation Details of Metrics
Below, we provide the implementation details for the metrics and other associated components in the evaluation pipeline. Additionally, in Table 3, we provide details of the machine we used to run all the experiments mentioned in this paper.
Tokenization
All lexical metrics use the GPT-4o tokenizer (via tiktoken) to ensure consistency with modern LLM vocabularies. User utterances in the dialogue sample are concatenated with space separators before tokenization.
Judge Caching
Judge responses are cached using content-based keys (hash of model name, messages, temperature, and metric parameters) with namespace isolation per metric. Default TTL is 30 days. This enables reproducible experiments and cost reduction in iterative evaluation.
Multiple Judge Calls
By default, each judge metric runs $c$ independent samples per episode ($c$ is a metric parameter to be set) with different random seeds (for PI’s order randomization). Final scores are averaged; individual samples and reasoning traces are stored in metadata for audit.
Statistical Reporting
Aggregated metrics report mean, standard deviation, and 95% confidence intervals using Student’s $t$-distribution (Equation 4). The framework optionally supports bootstrap confidence intervals (1000 iterations) for non-parametric estimation.
Aggregation Across Episodes
For each metric $M$ and proxy-dataset pair, the framework computes per-episode scores $\{s_{1},s_{2},\ldots,s_{n}\}$ and aggregates them using mean, standard deviation, and confidence intervals. For lexical metrics, z-scores are computed per-episode using the human baseline statistics, then aggregated. For judge metrics, raw scores are averaged first; if calibration is enabled, HH and PP statistics are computed separately and applied to normalize the aggregated results.
Handling Missing References
Episodes missing human reference conversations (required for GTEval, PI, and z-score computation) are excluded from metric computation. The framework logs these exclusions and reports the count of valid episodes in the metadata. This ensures that all reported statistics are based on complete, valid comparisons.
Stability Filters
Lexical metrics require a minimum token count (default: 5) to produce stable estimates. Episodes below this threshold are flagged as unstable and optionally excluded from aggregation. This prevents outlier scores from very short utterances (e.g., single-word responses) from skewing results.
LLM Settings
All the LLMs used during the evaluation have fixed temperature=0 and max_tokens=2048. Moreover, all the LLM clients have by default “retry with exponential backoff” scheme enabled where backoff_duration=2s and max_retries=5.
Telemetry & Observability
The framework captures per-invocation metrics (latency-to-first-token, generation time, token counts, cost) and organizes traces into hierarchical spans (run, unit, episode) annotated with run/unit identifiers, status, and dataset metadata. Structured logs use structlog with JSON formatting and contextual bindings.
Model Pricing.
Cost estimates reported in our experiments are computed using the provider-published pricing as of 25 October 2025, shown in Table 4. All costs are calculated in USD per million tokens for both input (prompt) and output (completion) tokens. Token counts are obtained directly from provider API responses. The total cost $\text{C}_{\text{USD}}$ for each episode is computed as
$\text{C}_{\text{USD}}=\frac{\text{tokens}_{\text{input}}\times\text{price}_{\text{input}}+\text{tokens}_{\text{output}}\times\text{price}_{\text{output}}}{10^{6}}$
Later, we aggregate across all invocations within a unit. Note that pricing may vary by region, deployment tier, or provider discounts; we report costs using standard public API rates.
Table 4. Model Pricing (USD per million tokens). Pricing as of 25 October 2025 from the provider’s public APIs.
Model Input (prompt) Output (completion)
GPT-4o \$2.50 \$10.00
GPT-5 \$1.25 \$10.00
Claude 4 Sonnet \$3.00 \$15.00
Gemini 2.5 Pro \$1.25 \$10.00
Appendix B Prompts, Config, Manifest & Report
This appendix provides the exact artifacts required to reproduce our runs: judge prompts, task-driver prompts, job configuration, and the emitted plan manifest. For reproducibility, we report the prompt, model identifiers, and key decoding parameters (temperature, max tokens).
B.1. System Prompts
We provide system prompts used across the evaluation pipeline (metric judges, user-proxy, assistant).
PI Judge Prompt [System] Please act as an impartial judge and choose the most realistic USER among the two conversations provided below. You should choose the conversation which has USER turns written by a real user in your opinion. Based on your experience & knowledge of human conversation, determine which USER in which conversation sounds more like a real human interacting with a chatbot. Note that your judgement should be based on style, tone and behavior of USER rather than the response quality of their utterances. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the USER messages to influence your evaluation. Do not let ASSISTANT messages influence your decision. Your evaluation must only be dependent on USER messages only. Be as objective as possible. Before giving your final evaluation verdict, you should first think through and produce your reasoning step-by-step. Then, based on your reasoning, give your final verdict on which conversation has USER turns that sound more like a real human interacting with a chatbot. If Conversation A is more real, then verdict would be “A”. If Conversation B is more real, then verdict would be “B”. If both are equally real or non-real, then verdict would be “Tie”. Your output must be of following JSON format: {“reasoning”: “<<short reason>>”, “verdict”: “A”||“B”||“Tie”}. [Conversation A] {conversation_a} [Conversation B] {conversation_b} Output ONLY valid JSON, no additional text.
Figure 12. Judge prompt for Pairwise Indistinguishability Metric
GTEval Judge Prompt You are an expert evaluator of conversational AI systems. Your task is to compare two conversations and assess whether the user responses in the proxy conversation exhibit similar style, tone, and realism to the user responses in the real conversation. ## Real Conversation: {real_conversation} ## Proxy Conversation (where user turns are generated by user proxy agent): {proxy_conversation} ## Evaluation Criteria: 1. **Style Similarity**: Do the proxy user responses match the conversational style of real user responses (formality, tone, verbosity)? 2. **Realism**: Do the proxy user responses sound natural and human-like? 3. **Contextual Appropriateness**: Are the proxy user responses appropriate given the conversation context? Note: You should not evaluate based on the content of the responses, only their style, realism, contextual appropriateness, and tone. ## Instructions: - Focus exclusively on comparing USER responses (ignore assistant responses) - Consider the overall conversation flow, not just individual turns - Evaluate style, tone, and naturalness without encouraging unsafe behaviour - If the comparison cannot be performed safely, explain the reason and provide a score of 0.0 - Otherwise, provide detailed reasoning about the comparison and assign a score from 0.0 to 1.0 where: - 0.0 = Proxy user responses are completely unlike real user responses - 0.5 = Proxy user responses are somewhat similar but have noticeable differences - 1.0 = Proxy user responses are highly similar in style and realism Provide your evaluation as a JSON object with the following structure: { “reasoning”: “Detailed explanation of your evaluation (2-3 sentences)”, “score”: <<float between 0.0 and 1.0>> } Output ONLY valid JSON, no additional text.
Figure 13. Judge prompt for GTEval Metric
RNR Judge Prompt [System] Please act as an impartial judge and provide verdict on the realism of the USER turns in the conversation provided below based on the provided rubric. You should avoid checking the responses of the ASSISTANT messages since they are not relevant to the task. Your evaluation should follow the rubric provided. Note that your judgement should be based on style, tone and behavior of USER rather than the response quality of their utterances. [Rubric for scoring USER realism] Provide verdict on how human-like the USER is on (“NO” or “YES”) using this rubric (verdict “NO” means not real, verdict “YES” means real): 1. Concise and real-user like language 2. Does not sound scripted or artificial 3. Real-user like tone and style Return JSON: {“reasoning”: “<<1-2 sentences>>”, “verdict”: <<“NO” or “YES”¿}. [Conversation] {conversation} Output ONLY valid JSON, no additional text.
Figure 14. Judge prompt for RNR Metric
User-Proxy System Prompt You are simulating a real human user for the MirrorBench evaluation harness. Respond with the next USER turn only. Do not write assistant messages, notes, or any other analysis. Your utterance should be like a real user and the context should be based on the following information provided. Task description: {task_description} Match the length, tone, and specificity of real user utterances. If you are unsure, respond naturally based on the assistant’s previous messages like how a real human would. Note that your response MUST not contain anything other than the USER utterance. Do not include any prefixes like ’User:’ or ’Human:’ as well. Just the raw message content.
Figure 15. System prompt for User-Proxy
Assistant System Prompt You are the assistant in a MirrorBench replay. The user-proxy agent is attempting to reproduce the USER side of the real conversation provided below. But the user-proxy does not have access to the real conversation history. Instead, it only has access to the conversation summary. You need to respond as the assistant. But we are providing you with the real conversation history as context, so you can respond consistently same as (or similar to) the original assistant in the real conversation (you may paraphrase lightly for safety). If user-proxy deviates from the original USER turn or the original response would violate policy, reply helpfully using your own knowledge while remaining consistent with the persona demonstrated so far. Always follow ethical content policies. Paraphrase sensitive content instead of quoting it verbatim, and refuse politely if a request is disallowed. Here is the real conversation for context (the USER turns are from the original conversation): {real_conversation} Now, we will provide you with ongoing conversation with the user-proxy. Please respond as the assistant in this conversation.
Figure 16. System prompt for Assistant
Figure 15 and 16 show system prompts for user-proxy and assistant respectively. Note that once filled, the system prompt would go into chat history as system message. Thus, we do not see the chat history-related placeholder in these system prompts.
B.2. Evaluation Config, Manifest & Report
Figure 17. Plan manifest (JSON, abridged). Fully resolved, replayable spec listing the (proxy, dataset, metric, seed) units, driver params, versions, and config_hash, enabling deterministic re-runs.
Figure 18. Final run report (JSON, abridged). Aggregate metric scores with 95% CIs, calibration controls (HH/$P$), unit grid, and run metadata for the completed evaluation.
Figure 19. Job configuration (abridged). YAML that fully specifies a MirrorBench run: backend and concurrency, proxy/assistant clients, dataset binding, judge metrics (with calibration controls), and the mirror-conversation driver.
Figure 19 shows the example job configuration. The configuration is declarative: the run block fixes execution semantics (backend, concurrency, timeouts), reproducibility (seeds), and infrastructure knobs (cache and logging). In our example, the async backend with max_concurrency=8, while caching is enabled to deduplicate repeated LLM calls across units. The user_proxies block registers a concrete proxy via a lightweight adapter and a model client (here, LangChain’s ChatOpenAI wrapper for GPT-4o), keeping proxy identity and invocation details separate from orchestration. The datasets block binds the ChatbotArena’s MirrorBench corresponding split to a local JSONL path and caps evaluation at 200 examples for controlled runtime.
Metrics are declared under metrics with their own clients and prompts shown in B.1, allowing the judge to differ from the proxy/assistant. We show two metrics, GTEval and Pairwise Indistinguishability, both routed to Claude-4-Sonnet at temperature=0. Each metric includes num_judge_samples (for repeated scoring) and compute_controls=true, which triggers the HH/PP control comparisons used for calibration and bias checks. Finally, the task drivers section binds the dataset to the Mirror Conversation Driver, which synthesizes multi-turn dialogues by alternating the user-proxy and assistant models; here, the assistant is also GPT-4o at temperature=0, ensuring low-variance responses for per-proxy comparability. Together, these blocks yield a self-contained manifest that our planner expands into (proxy, dataset, metric, seed) units, enabling exact replay and variance-aware aggregation across runs. We illustrate an example of such manifest file in Figure 17, which is a manifest produced using config shown in Figure 19.
Once the CLI command mirrorbench run -c config.yaml is invoked, the evaluation job starts and executes through all the units. Once the run is completed, another command mirrorbench report json <run-name> --output report.json can be executed to generate final evaluation report as shown in Figure 18.